PET has a Young's, Modulus of 3.5 GPa and PLGA 1.2 GPa. A polymer blend is prepared with 50:50 mixture (PET:PlGA) . What is the Young's, Modulus of this blend (use mixture rule). Calculate the percent load borne by PET when the blend is subjected to a load. If after 10 days, half of PET and one third of PLGA had degraded, what is the Young's Modulus of this blend.
1
Expert's answer
2020-04-14T01:44:55-0400

"Young's modulus=Y_{PET}=" 3.5 GPa, "Y_{GPA}= 1.2 GPa"

It has mixture of 50:50

"Y_m= \frac {Y_{PET}}{2} + \frac {Y_{GPA}}{2}"


"Y_m= \frac {3.5}{2} + \frac {1.2}{2}"


Y"_m= 1.75+.6= 2.35 GPa"


Let . P be the load applied on blend mixture


Total elongation = elongation of 1 + elongation of 2

so here Youngs modulues for mixture after 10 days as per rule of degradation



"Ym =" "\frac{3.5}{2} +\frac{1.2}{3}= 1.74 + .4= 2.14" GPa


This Youngs modulus after 10 days